An air column of variable length is open at one end and closed at the other end. If it has a 5th resonant length of 1.25 m,

a) Calculate the wavelength of the sound wave.

b) Calculate the 2nd resonant length of the air column.


1
Expert's answer
2021-02-01T12:41:25-0500

a)


"\lambda=\frac{4}{5}L\\\lambda=\frac{4}{5}1.25=1\ m"

b)


"L_3=\frac{3}{4}\lambda=0.75\ m"